if an object of mass 15kg is travelling at 5m/s, how much kinetic energy does it have?


1
Expert's answer
2021-03-08T08:30:50-0500
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot15\ kg\cdot(5\ \dfrac{m}{s})^2=187.5\ J."